# Back to the Color: Learning Depth to Specific Color Transformation for Unsupervised Depth Estimation

### v2

@@ Abstract
Virtual engines have the capability to generate dense depth maps for various synthetic scenes, making them invaluable for training depth estimation models. However, synthetic colors often exhibit significant discrepancies compared to real-world colors, thereby posing challenges for depth estimation in real-world scenes, particularly in complex and uncertain environments encountered in unsupervised monocular depth estimation tasks. To address this issue, we propose Back2Color, a framework that predicts realistic colors from depth utilizing a model trained on real-world data, thus facilitating the transformation of synthetic colors into real-world counterparts. Additionally, by employing the Syn-Real CutMix method for joint training with both real-world unsupervised and synthetic supervised depth samples, we achieve improved performance in monocular depth estimation for real-world scenes. Moreover, to comprehensively address the impact of non-rigid motions on depth estimation, we propose an auto-learning uncertainty temporal-spatial fusion method (Auto-UTSF), which integrates the benefits of unsupervised learning in both temporal and spatial dimensions. Furthermore, we design a depth estimation network (VADepth) based on the Vision Attention Network. Our Back2Color framework demonstrates state-of-the-art performance, as evidenced by improvements in performance metrics and the production of fine-grained details in our predictions, particularly on challenging datasets such as Cityscapes for unsupervised depth estimation. Code is available at https://github.com/BlueEg/back2color/.
@@
Back to the Color, auto-learning uncertainty, Vision Attention Depth NetworkI Introduction
@@
In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[1]. Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[2, 3]. Even with the refinement of advanced LiDAR systems[3], their data are inherently sparse, comprising only discrete depth points. Even the most precise scanning LiDAR systems can provide only sparse depth points. Unsupervised depth estimation, powered by epipolar geometry[4] using purely visual approaches does not require additional annotation for depth information other than camera parameters. Unsupervised depth estimation eliminates the need to create specific datasets for particular scenes[5]. By utilizing binocular and motion disparity clues, it mimics the structure of human vision, making it efficient and cost-effective[6]. This is why many researchers and companies are actively pursuing this approach[7]. Monodepth2[8], the most renowned framework for unsupervised monocular depth estimation, has established itself as a classic and successful method. By reprojecting related frames(temporal and spatial) and regressing depth values. This adaptability enables Monodepth2 to obtain depth information for various scenes through adaptive fine-tuning, making it highly versatile. However, despite its strengths, Monodepth2 has several limitations. Firstly, the auto-masking for handling non-rigid moving objects, follows a simple rule: the pixel error after reprojection should be less than before reprojection. This pixel-level determination cannot ensure effectiveness in smooth regions. Additionally, when applied to more complex datasets such as Cityscapes[9], it almost fails. Secondly, photometric reprojection error also fails in areas with no texture or with texture-less background objects, leading to depth prediction failure. Lastly, methods based on photometric reprojection matching lack prior knowledge of the objective structure. These issues make it challenging for unsupervised depth estimation methods to be applied in real world.
@@
With the continuous advancement of computer graphics and virtual engine technologies, significant progress has been made in the creation and application of virtual urban scene datasets[10, 11, 9, 12, 13]. The evolution of 3D virtual engines, such as Unity and Unreal Engine[14, 15], has played a crucial role in enhancing the realism and visualization effects of virtual urban scenes. These engines not only facilitate finer scene modeling but also provide interactivity and dynamism, making virtual urban scene data more closely resemble the real world. Synthetic datasets [10] can provide precise qualitative data for virtual urban scenes, including depth, semantic labels, and more. Therefore, it would be desirable if synthetic depth could be used to assist in the unsupervised depth estimation of real-world [16]. Compared to synthetic datasets, real-world depth estimation datasets[17, 9] have the following notable drawbacks: firstly, synthetic depth data is dense, whereas depth data collected by LiDAR is sparse and discrete[17]. Secondly, depth computed through stereo matching suffers from occlusions and lacks accuracy due to the short baseline[9]. Additionally, sparse depth sampling does not provide sufficient information for learning when obstacles are small or distant (resulting in small apparent areas), making it challenging for the network to accurately reconstruct their shapes.
@@
In this paper, our Back2Color learns the mapping process in the target domain (Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains. Additionally, we introduced an improved version of CutMix[18], termed Syn-Real CutMix, to effectively combine real-world, synthesized, and syn-to-real data in both batch and image spatial dimensions. By incorporating Syn-Real CutMix into our training process, we were able to enrich the training dataset with more diverse color-to-depth knowledge extracted from synthetic scenes. Furthermore, the joint training with diverse data sources enables the model to capture a wider range of color and depth variations present in real-world scenes, thus enhancing its robustness and generalization capability.
Additionally, we addressed the issue of chaotic depth clouds(5) by proposing a Background Textureless regularization loss (sky regularization loss), aiming to narrow the gap between predicted depths from real-world and synthesized data. Through joint training, our method enables the unsupervised monocular depth estimation to leverage knowledge from synthesized data, resulting in the generation of dense and sharp depths.
@@
We designed an efficient and lightweight Depth Net based on the Vision Attention Network. This network features fewer parameters and lower computational complexity compared to transformer-based networks, while achieving superior performance.
Furthermore, to more robustly recognize non-rigid moving objects and address occlusion issues in stereo systems, we enhance the method based on Auto-Uncertainty by introducing Auto-Spatial-Temporal Uncertainty Fusion, which leverages complementary spatial and temporal advantages in the unsupervised depth estimation process.
@@ II-A Unsupervised Depth Estimation
Unsupervised depth estimation aims to predict distances from stereo views or video sequences. Initial methods like those by Godard et al. [21] introduced disparity consistency loss, while Zhou et al. [22] pioneered depth and pose estimation simultaneously. Recent advancements [23, 24, 25, 26, 27] refine network structures for higher accuracy. Techniques like PackNet [23] and CADepth [24] enhance geometric and appearance information propagation. HR-Depth [25] and DIFFNet [26] utilize HRNet for better resolution utilization. RA-Depth [27] employs HRNet for both encoding and decoding. In natural language processing, the Transformer architecture [28] has been successful, later adapted to computer vision tasks [29, 30]. It serves as the backbone for depth estimation [31, 32, 33, 34, 35], but faces challenges like computational demands and real-time constraints. Existing methods prioritize model architecture, often overlooking crucial obstacles and targets.
@@ II-C Style transfer in urban scenes
Research on style transfer for urban scene images has long been ongoing due to its importance in autonomous driving technology. Isola et al. [39] proposed a conditional GAN framework for generating urban scene images from semantic segmentation images of Cityscapes in 2017. However, generative models may produce random and plausible objects, disrupting the scene structure, as shown in their experiments. Atapour et al. [40] utilized style transfer and adversarial training to convert real-world color images into synthetic environment-style images. Huang et al. [41] applied GAN-based domain transfer for data augmentation in urban scene weather-time conversion. To minimize the visual gap between source and target domains, Gomez et al. [42] aligned colors across different domains in the Lab color space. Zheng et al. [43] used GANs to convert synthetic colors to real-world colors.
@@ III Methodology
In this section, we will first introduce our Back2Color method, which involves learning the depth-to-color mapping relationship in the target domain to train the Color Net. This method also includes the spatial-batch joint training technique named Real-Syn CutMix and an approach for handling background textureless objects. Next, We will introduce our Depth Net based on the Vision Attention Network, which balances performance and lightweight design. Finally, we will present a more robust method, Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF), to recognize and solve non-rigid motion problems.
@@ III-A1 Transform depth to specific colors
the significant gap between synthetic and real-world color images arises from the inherent complexity of the real world. Colors in real-world scenes are influenced by various factors, including lighting intensity, surface material of obstacles, and smoothness. Each of these factors is intricate and highly variable, making them challenging to simulate realistically using typical 3D engines. As a result, discerning the origin of an image intuitively is feasible, both for humans and deep learning alike. Consequently, directly combining real-world and synthetic datasets for training essentially involves dealing with two distinct domains. Our Back2Color approach learns the mapping relationship in the target domain(Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains.
Predicting depth from color entails extrapolating 3D information from optical data, mirroring the process carried out by the human brain. Throughout training, neural networks assimilate various factors such as obstacle features, perspective and occlusion relationships, fine-grained surface details, apparent surface areas, as well as lighting and shadow effects[44]. It acquired features are subsequently leveraged to estimate the depth of obstacles.
@@
And utilizing the input of the Depth Net for Sup:
Here, represents the weights that store specific dataset color mapping information, such as Kitti and Cityscapes. Fortunately, we found that this prediction is feasible, neural networks are able to capture color information under specific target scenes. Next, we need to investigate whether we can employ to convert predicted in synthesized datasets into color that resemble the Kitti() color style: Using the specific Color Net to colorize predicted synthetic depth, as shown in Fig.3:
@@ III-A2 Syn-Real CutMix
the renowned augmentation technique CutMix[18], replaces removed pixels with patches from other images rather than simply removing them. This strategy eliminates uninformative pixels during training, enhancing efficiency, while maintaining the advantages of regional dropout to focus on non-discriminative obstacle parts. Additionally, introducing extra patches improves the model’s localization ability, requiring identification of obstacles from partial views, In our work, we apply this technique to jointly train both real-world and synthetic datasets, named Syn-Real CutMix.
The goal of CutMix in our work is to generate a new training pair . Unlike CutMix[18], our samples are not two unrelated samples; instead, they correspond to different colored images of the same depth label . Therefore, we only need to merge color images without combining new labels for them. To increase the randomness of training samples. As shown in Fig.4, given image samples as , we firstly randomly swap . Then, randomly crop a block from with a height and width greater than half of the original image, and replace it with the corresponding position in :
@@
Now, we combine Syn-Real enhanced images and Real world target image into one batch for training. Then, we can train monocular depth estimation networks simultaneously on real-world, synthetic, and Syn-to-Real images within a batch. Spatial randomization also encourages the network to focus more on the image structure rather than its specific colors.
@@ III-A3 Background textureless regularization loss
Here, is inferred from a pre-trained semantic segmentation network[48]. We also have an intuitive prior that the sky should only exist above in the view, so we added a decay coefficient that decreases with the image height:
@@ III-B Vision Attention Depth Net
Every high-performing depth estimation network requires an effective and robust backbone to extract hierarchical features from natural images. Commonly used backbones include ResNet[49] (Monodepth2[8], HRDepth[25], CADepth[24]), HRNet[50] (DIFFNet[26], RA-Depth[27]), and Transformer-based backbones like ViT[29], MPViT[35](MonoViT[33], TUSDepth[20]). With the evolution from ResNet to HRNet and then to Transformer-based approaches, the performance of depth estimation continues to improve, but accompanied by increasing computational complexity and params. We consistently believe that depth estimation tasks, especially those based on street scenes, need to consider the requirements of their usage scenarios (real-time, lightweight). Hence, we aim to find a more efficient and effective backbone than Transformer methods.
@@
Due to its outstanding performance and efficiency, we have introduced the VAN model into unsupervised monocular depth estimation tasks, and adapt to some fine-tuning. The most significant change is that, to reduce computational complexity, VAN downsamples the input resolution to during the first stem stage. While this downsampling is acceptable for classification tasks, it poses challenges for high-resolution recovery in image regression tasks. Therefore, without altering the original VAN structure, we have added an additional stem head to acquire features at resolution, as shown in Fig.6. After extracting features from various VAN stages, we equipped them with a high-resolution depth decoder[27] to output the final depth. During implementation, we observed no difference between outputting a single-size depth image and outputting multi-scale depth images. Therefore, our network only outputs depth maps that are the same size as the input images.
@@ III-C Auto-learning uncertainty temporal-spatial fusion
When testing our designed VADepth model, we found that the Auto-masking method almost failed, as shown in Fig.7. This issue is more severe in the Cityscapes[9]. Upon inspecting the output of Auto-masking, we discovered that it couldn’t perfectly shield non-rigid objects. We attribute this to the model’s robustness, the limitations of the Auto-masking method, and the complexity of the Cityscapes[9, 51].
@@
To combine photometric loss across different frames, they specify . Hence, they combine the output uncertainty map with softmax: , to ensure that . The final joint optimization function is written as:
Using this formula, the network will automatically learn the uncertain regions within the loss, forming the final weights for multi-frame loss through comparison (areas where the loss cannot be reduced through optimization). In the temporal multi-frame reprojection loss combination, high-uncertainty regions are the occluded areas revealed due to reprojection. Meanwhile, in the process of spatial and temporal reprojection loss combination, high-uncertainty regions should be the non-rigid motion areas that temporal reprojection cannot resolve.
and represent the temporal domain and the spatial domain, respectively. and represent the temporal-domain uncertainty and the spatial-domain uncertainty, , . represents the source frame in a stereo setup. Since the Kitti dataset randomly selects either the left or right camera image as the target frame, the other frame becomes the source frame.
Regarding how to obtain temporal and spatial uncertainty maps, we consider referencing the method used to generate temporal uncertainty maps by inputting related frames containing uncertainty information. For spatial reprojection, we can input and , but there is no single related frame that uniquely represents temporal reprojection. We attempt to input and as representatives. By comparing the output uncertainty maps when calculating the loss, we can form reasonable uncertainty maps, as shown in Fig. 9Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues. We improved and applied this method to fuse temporal and spatial reprojection errors. Since non-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty. Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process. By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depth images.
Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues[20]. We improved and applied this method to fuse temporal and spatial reprojection losses. Since no-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty, as shown in Fig.10. Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process. By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depth images, as shown in Fig.9.

### v3

@@ Abstract
Virtual engines can generate dense depth maps for various synthetic scenes, making them invaluable for training depth estimation models. However, discrepancies between synthetic and real-world colors pose significant challenges for depth estimation in real-world scenes, especially in complex and uncertain environments encountered in unsupervised monocular depth estimation tasks. To address this issue, we propose Back2Color, a framework that predicts realistic colors from depth using a model trained on real-world data, thus transforming synthetic colors into their real-world counterparts. Additionally, we introduce the Syn-Real CutMix method for joint training with both real-world unsupervised and synthetic supervised depth samples, enhancing monocular depth estimation performance in real-world scenes. Furthermore, to mitigate the impact of non-rigid motions on depth estimation, we present an auto-learning uncertainty temporal-spatial fusion method (Auto-UTSF), which leverages the strengths of unsupervised learning in both temporal and spatial dimensions. We also designed VADepth, based on the Vision Attention Network, which offers lower computational complexity and higher accuracy than transformers. Our Back2Color framework achieves state-of-the-art performance on the Kitti dataset, as evidenced by improvements in performance metrics and the production of fine-grained details. This is particularly evident on more challenging datasets such as Cityscapes for unsupervised depth estimation. Code is available at https://github.com/BlueEg/back2color/.
@@
Back2Color, Syn-Real CutMix, auto-learning uncertainty, Vision Attention Depth Network
I Introduction
@@
In the real world, even after prolonged evolution, humans are still limited to qualitatively analyzing the depth of specific points in their field of view to assist in their actions[3]. Despite significant technological advances, the acquisition of precise and dense depth information remains a formidable challenge[4, 5]. Even with the refinement of advanced LiDAR systems[5], their data are inherently sparse, comprising only discrete depth points. Unsupervised depth estimation(UDE), powered by epipolar geometry[6] using purely visual approaches does not require additional annotation for depth information other than camera parameters. UDE eliminates the necessity of annotating depth labels for specific scenes[7]. By utilizing binocular and motion disparity clues, it mimics the structure of human vision, making it efficient and cost-effective[8]. This is why many researchers and companies are actively pursuing this approach[9]. Monodepth2[10], the most renowned framework for unsupervised monocular depth estimation, has established itself as a classic and successful method. By reprojecting related frames(temporal and spatial) and regressing depth values. This adaptability enables Monodepth2 to obtain depth information for various scenes through adaptive fine-tuning, making it highly versatile. However, despite its strengths, Monodepth2 has several limitations. Firstly, the auto-masking for handling non-rigid moving objects, follows a simple rule: the pixel error after reprojection should be less than before reprojection. This pixel-level determination cannot actively detect motion[11] and cannot ensure effectiveness in smooth regions. Additionally, when applied to more complex datasets such as Cityscapes[12], it almost fails. Secondly, photometric reprojection error also fails in areas with no texture or with texture-less background objects, leading to depth prediction failure. Lastly, methods based on photometric reprojection matching lack prior knowledge of the objective structure. These issues make it challenging for UDE methods to be applied in real world.
@@
With the continuous advancement of computer graphics and virtual engine technologies, significant progress has been made in the creation and application of virtual urban scene datasets[13, 14, 12, 15, 16]. The evolution of 3D virtual engines, such as Unity and Unreal Engine[17, 18], has played a crucial role in enhancing the realism and visualization effects of virtual urban scenes. These engines not only facilitate finer scene modeling but also provide interactivity and dynamism, making virtual urban scene data more closely resemble the real world. Synthetic datasets can provide precise qualitative data for virtual urban scenes, including depth, semantic labels, and more. Therefore, it would be desirable if various synthetic scenes and dense depth could be used to assist in the UDE of the real-world[19]. Compared to synthetic datasets, real-world depth estimation datasets[20, 12] have the following notable drawbacks: firstly, synthetic depth data is dense, whereas depth data collected by LiDAR is sparse and discrete[20]. Secondly, depth computed through stereo matching suffers from occlusions and lacks accuracy due to the short baseline[12]. Additionally, sparse depth sampling does not provide sufficient information for learning when obstacles are small or distant (resulting in small apparent areas), making it challenging for the network to accurately reconstruct their shapes.
@@
In this paper, our Back2Color learns the mapping process in the target domain (Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains. Additionally, we introduced an improved version of CutMix[21], termed Syn-Real CutMix, to effectively combine real-world, synthesized, and syn-to-real data in both batch and image spatial dimensions. By incorporating Syn-Real CutMix into our training process, we were able to enrich the training dataset with more diverse color-to-depth knowledge extracted from synthetic scenes. Furthermore, the joint training with diverse data sources enables the model to capture a wider range of color and depth variations present in real-world scenes, thus enhancing its robustness and generalization capability. Additionally, aiming to narrow the gap between predicted depths from real-world and synthesized depth, we addressed the issue of chaotic depth clouds (5) by introducing a Background Textureless Regularization Loss, also known as Sky Regularization Loss. Through joint training, our method enables the unsupervised monocular depth estimation to leverage knowledge from synthesized data, resulting in the generation of dense and sharp depths.
@@
We designed an efficient and lightweight Depth Net based on the Vision Attention Network. This network boasts fewer parameters and lower computational complexity compared to transformer-based networks, while delivering superior performance.
To more robustly recognize non-rigid moving objects and address occlusion issues in stereo systems, we enhance the method based on Auto-Uncertainty by introducing Auto-Spatial-Temporal Uncertainty Fusion, which leverages complementary spatial and temporal advantages in the unsupervised depth estimation process.
@@ II-A Unsupervised Depth Estimation
Unsupervised depth estimation aims to predict distances from stereo views or video sequences. Initial methods like those by Godard et al.[27] introduced disparity consistency loss, while Zhou et al.[28] pioneered depth and pose estimation simultaneously. Recent advancements[29, 30, 31, 32, 33] refine network structures for higher accuracy. Techniques like PackNet[29] and CADepth[30] enhance geometric and appearance information propagation. HR-Depth[31] and DIFFNet[32] utilize HRNet for better resolution utilization. RA-Depth[33] employs HRNet for both encoding and decoding. In natural language processing, the Transformer architecture[34] has been successful, later adapted to computer vision tasks[35, 36]. It serves as the backbone for depth estimation[37, 38, 2, 1, 39], but faces challenges like computational demands and real-time constraints. Existing methods prioritize model architecture, often overlooking crucial obstacles and targets. Additionally, depth estimation models also face the issue of overfitting on specific datasets and a lack of continuous learning capability[40].
@@ II-C Style transfer in urban scenes
Research on style transfer for urban scene images has long been ongoing due to its importance in autonomous driving technology. Isola et al.[44] proposed a conditional GAN framework for generating urban scene images from semantic segmentation images of Cityscapes in 2017. However, generative models may produce random and plausible objects, disrupting the scene structure, as shown in their experiments. Atapour et al.[45] utilized style transfer and adversarial training to convert real-world color images into synthetic environment-style images. Huang et al.[46] applied GAN-based domain transfer for data augmentation in urban scene weather-time conversion. To minimize the gap between source and target domains, Gomez et al.[47] aligned colors across different domains in the Lab color space. Zheng et al.[48] used GANs to convert synthetic colors to real-world colors.
@@ III Methodology
In this section, we introduce the Back2Color framework, which masters the depth-to-color mapping within the target domain to train the Color Net (III-A1). Back2Color is enhanced by Real-Syn CutMix (III-A2) and a strategy for handling textureless background objects (III-A3). We then discuss the VADepth, based on the Vision Attention Network, designed for high performance and lightweight architecture (III-B). Lastly, we describe the Auto-learning Uncertainty Temporal-Spatial Fusion (Auto-UTSF) (III-C) for addressing non-rigid motion challenges. In this paper, Depth and Disparity are equivalent.
@@ III-A1 Transform depth to specific colors
the significant gap between synthetic and real-world color images arises from the inherent complexity of the real world. Colors in real-world scenes are influenced by various factors, including lighting intensity, surface material of obstacles, and smoothness, and can also be significantly degraded[49]. Each of these factors is intricate and highly variable, making them challenging to simulate realistically using typical 3D engines. As a result, discerning the origin of an image intuitively is easy, both for humans and deep learning alike. Consequently, directly combining real-world and synthetic datasets for training essentially involves dealing with the integration issues of two distinct domains. Our Back2Color approach learns the mapping relationship in the target domain(Real). This mapping is then applied to the source domain (Synthetic) to reduce the gaps between colors of the source and target domains.
@@
and utilizing the input of the Depth Net for Supervision:
where represents the weights that store specific dataset color mapping information, such as Kitti and Cityscapes. Fortunately, we found that this prediction is feasible, neural networks are able to capture color information under specific target scenes. Next, we need to investigate whether we can employ to convert predicted in synthesized datasets into color that resemble the Kitti() color style: Using the specific Color Net to colorize predicted synthetic depth, as shown in Fig.3:
Our experiments have shown that by referencing color information from the target domain, we can learn its depth-to-color mapping and accurately predict color images in the source domain that closely resemble those in the target domain.
@@ III-A2 Syn-Real CutMix
CutMix[21] replaces removed pixels with patches from other images rather than simply removing them. This strategy eliminates uninformative pixels during training, enhancing efficiency, while maintaining the advantages of regional dropout to focus on non-discriminative obstacle parts. Additionally, introducing extra patches improves the model’s localization ability, requiring identification of obstacles from partial views. In our work, we draw inspiration from the CutMix method to achieve different objectives and effects. We aim to dynamically fuse Syn2Real Colors and Synthetic Colors within a single batch for supervised training, a technique we have named Syn-Real CutMix.
The goal of CutMix in our work is to generate a new training pair . Unlike CutMix[21], our samples are not two unrelated samples; instead, they correspond to different colored images of the same depth label . Therefore, we only merge color images without combining new labels. To increase the randomness of training samples. As shown in Fig.4, given image samples as , we firstly randomly swap . Then, randomly crop a block from with a height and width greater than half of the original image, and replace it with the corresponding position in :
@@
Now, we combine Syn-Real enhanced images and Real world target image I into one batch for training. Then, we can train monocular depth estimation networks simultaneously on real-world, synthetic, and Syn-to-Real images within a batch. Spatial randomization also encourages the network to focus more on the image structure rather than its specific colors.
@@ III-A3 Background textureless regularization loss
where is inferred from a pre-trained semantic segmentation network[50]. We also have an intuitive prior that the sky should only exist above in the view, so we added a decay coefficient d that decreases with the image height:
@@ III-B Vision Attention Depth Net
Every high-performing depth estimation network requires an effective and robust backbone to extract hierarchical features from natural images. Commonly used backbones include ResNet[51] (Monodepth2[10], HRDepth[31], CADepth[30]), HRNet[52] (DIFFNet[32], RA-Depth[33]), and Transformer-based backbones like ViT[35], MPViT[39](TUSDepth[23], MonoViT[2]). With the evolution from ResNet to HRNet and then to Transformer-based approaches, the performance of depth estimation continues to improve, but accompanied by increasing computational complexity and params. We consistently believe that depth estimation tasks, especially those based on street scenes, need to consider the requirements of their usage scenarios (real-time, lightweight). Hence, we aim to find a more efficient and effective backbone than Transformer methods.
@@
Due to its outstanding performance and efficiency, we have introduced the VAN model into unsupervised monocular depth estimation tasks, and adapt to some fine-tuning. The most significant change is that, to reduce computational complexity, VAN downsamples the input resolution to during the first stem stage. While this downsampling is acceptable for classification tasks, it poses challenges for high-resolution recovery in image regression tasks. Therefore, without altering the original VAN structure, we have added an additional stem head to acquire features at resolution, as shown in Fig.6. After extracting features from various VAN stages, we equipped them with a high-resolution depth decoder[33] to output the final depth. During implementation, we observed no difference between outputting a single-size depth image and outputting multi-scale depth images. Therefore, our network only outputs depth maps with the same size as the input.
@@ III-C Auto-learning uncertainty temporal-spatial fusion
When testing our designed VADepth model, we found that the Auto-masking[10] method almost failed, as shown in Fig.7. This issue is more severe in the Cityscapes[12]. Upon inspecting the output of Auto-masking, we discovered that it couldn’t perfectly shield non-rigid objects. We attribute this to the model’s robustness, the limitations of the Auto-masking method, and the complexity of the Cityscapes[12, 53].
@@
To combine photometric loss across different frames, Zhu et al.[54] specify . Hence, they combine the output uncertainty map with softmax: , to ensure that . The final joint optimization function is written as:
Using this formula, the network will automatically learn the uncertain regions(areas where the loss cannot be reduced through optimization) within the loss, forming the final weights for multi-frame loss through comparison. In the temporal multi-frame reprojection loss combination, high-uncertainty regions are the frames’ occluded areas revealed due to reprojection. Meanwhile, in the spatial and temporal reprojection loss combination, high-uncertainty regions in the temporal reprojection loss map are the non-rigid motion areas that can be accurately reprojected in the spatial.
where and represent the temporal domain and the spatial domain, respectively. and represent the temporal-domain uncertainty and the spatial-domain uncertainty, , . represents the source frame in a stereo setup. Since the Kitti dataset randomly selects either the left or right camera image as the target frame, the other frame becomes the source frame.
Regarding how to obtain temporal and spatial uncertainty maps, we consider referencing the method used to generate temporal uncertainty maps by inputting related frames containing uncertainty information. For spatial reprojection, we can input and , but there is no single related frame that uniquely represents temporal reprojection. We attempt to input and as representatives. By comparing the output uncertainty maps when calculating the loss, we can form reasonable uncertainty maps, as shown in Fig. 9. Applying uncertainty learning in temporal reprojection can address occlusion and field contraction issues. We improved and applied this method to fuse temporal and spatial reprojection errors. Since non-rigid objects are rigid in the spatial domain, the temporal domain generates significant uncertainty(Fig.10). Additionally, high uncertainty still arises from occluded regions during the spatial domain reprojection process. By fusing the uncertainties from both the temporal and spatial domains, we can ignore high-uncertainty areas in each domain when calculating the loss, leveraging the strengths of each domain to train better depths.